OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches two of quantity x^2-4/x-2 @agent0smith

OpenStudy (agent0smith):

First, try factoring the numerator.

OpenStudy (agent0smith):

I gotta go out for a bit, maybe @ranga can finish off.

OpenStudy (anonymous):

OKAY :(

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\frac{x^2-4}{x-2} } \] \[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\frac{(x-2)(x+2)}{x-2} } \] cancel out x-2 and plug in 2 for x.

OpenStudy (anonymous):

@agent0smith is it x+4

OpenStudy (anonymous):

wait nooo

OpenStudy (solomonzelman):

No

OpenStudy (anonymous):

i have (x-2)(x+2)/x-2

OpenStudy (ranga):

Solomon: Are you using Chrome? The math editor, draw, latex, none of them work in Firefox for me.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\frac{x^2-4}{x-2} } \] \[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\frac{(x-2)(x+2)}{x-2} } \] \[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\frac{\cancel{(x-2) }(x+2)}{\cancel{x-2 }} } \]

OpenStudy (ranga):

Solomon: All your replies are shown up raw.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~x+2}\] \[\huge\color{blue}{ \lim_{x \rightarrow 2}~~~\color{red} { 2 }+2}\] \[\huge\color{blue}{ \lim_{x \rightarrow 2}~~?}\] you tell me

OpenStudy (solomonzelman):

@melacho, do you see the huge blue fractions, or a bunch of not making sense of letters numbers and etc ?

OpenStudy (anonymous):

yes:/

OpenStudy (solomonzelman):

ranga I am using chrome, and equation editor; yes. this is not a yes or no question....

OpenStudy (anonymous):

do (x-2) and (x+2) cancel out?

OpenStudy (solomonzelman):

No, only x-2 and x-2, not x+2 and x-2

OpenStudy (anonymous):

so x+2

OpenStudy (solomonzelman):

yes, you would then have limit x--->2 x+2 so plug in 2 for x, and you get limit x--->2 ? tell me the answer,

OpenStudy (anonymous):

does not exist

OpenStudy (solomonzelman):

I am not going to continue helping you, if you don't know how to plug in 2 for x. you are solving a limit and at this stage you should know this.

OpenStudy (anonymous):

2^2-4/2-2=0/0

OpenStudy (solomonzelman):

no no no, 1) factor the top. rule to use: a^2-b^2=(a+b)(a-b) 2) simplify the fraction, by canceling out (x-2) on top and bottom 3) plug in 2 for x.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Yes, 4, good job!

OpenStudy (anonymous):

thanks c:

OpenStudy (solomonzelman):

yw